Suppose that diameters of a new species of apple have a bell-shaped distribution with a mean of 7.26 cm and a standard deviation of 0.43 cm. Using the empirical rule, what percentage of the apples have diameters that are less than 5.97 cm? Please do not round your answer. 
Answer #1

Answer:

Mean: 7.26cm

Standard deviation: 0.43cm

5.97= 7.26-3*0.43= mean -3*sd

By Empirical Rule, Approximately 99.7% of the data occurs within 3 SD’s of the mean.

Therefore 0.15% below mean -3*sd limit and 0.15% above mean +3*sd limit of data occurs.

Therefore the required percentage= 0.15%